Use the following quadratic equation to complete parts a and b below. You must show all work.
x2 + 4x - 21 = 0
a. Solve by factoring.
b. Solve using the quadratic formula.

Answer by stanbon(75887)

x^2 + 4x - 21 = 0
a. Solve by factoring.
x^2+7x-3x-21 = 0
x(x+7)-3(x+7) = 0
(x+7)(x-3) = 0
x = -7 or x = 3
=============================================
b. Solve using the quadratic formula.
x = [-4 +- sqrt(16 - 4*1*-21)]/2
---
x = [-4 +- sqrt(16+84)]/2
---
x = [-4 +- 10]/2
---
x = [-2+- 5]
---
x = -7 or x = 3
===================
Cheers,
Stan H.
===============